a sample contain N moles of diatomoc gas at temp T.molecules of gas gas get dissociated into atoms remaning temp constant.Find the change in the internal energy

Grade:12

1 Answers

AskiitianExpert Pramod-IIT-R
14 years ago
when gas is diatomic U1 = nCv * T = N * 5/2 * R * T when gas converted to mono atomic U2 = 2N * 3/2 * R * T U1-U2 = N * 5/2 * R * T - 2N * 3/2 * R * T = NRT (2.5-3.0) = -0.5NRT Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE & AIEEE preparation. All the best. Regards, Askiitians Experts Pramod Kumar IITR Alumni